### v1

@@ 4 Methodology
In this section, we give a detailed description of UACER whose overview is illustrated in Figure 1. We first introduce the diversified critic ensemble framework for robust adversarial RL and then develop the variance-based TDU Q-value aggregation function, followed by theoretical guarantees for TDU’s convergence properties and aggregation effectiveness. In principle, our method can be used in conjunction with most modern robust adversarial RL algorithms, including recent advances like QARL (Reddi et al., 2024) as well as classical approaches such as RARL (Pinto et al., 2017). For the sake of concreteness, we here provide the QARL-based implementation as our primary instantiation.
Figure 1: Schematic of the proposed UACER method.
4.1 Diversified Critic Ensemble
The Unified Framework.
@@ 4.1 Diversified Critic Ensemble
The UACER framework implements both protagonist and adversary as actor-critic agents based on the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018). Each agent consists of a single policy network, denoted as $\pi_{\phi}$, and an ensemble of $K$ soft Q-function networks $\{Q_{\theta_{k}}\}_{k=1}^{K}$. This design is motivated by the empirical insight, as shown in Figure LABEL:actor_ensemble, that the critic ensemble significantly enhances the robustness of Q-value estimation, especially under adversarial perturbations. By evaluating Q-values through multiple independent critics, the framework mitigates the influence of localized noise or perturbations on any single estimator, thereby yielding more stable and reliable value estimates. In contrast, although actor ensembles can improve sample efficiency and generalization (Lee et al., 2021; Yang et al., 2022), they often introduce additional training instability due to policy divergence among actors, which may produce conflicting behaviors that undermine the consistency and convergence of the learned policy.
@@
During the adversarial training phase, both agents interact synchronously with the environment through their current policies, generating transition tuples $\tau_{t}=(s_{t},a_{p_{t}},a_{a_{t}},r_{t},\\ s_{t+1})$ that are stored in the experience replay buffer $\mathcal{D}$. Given the symmetric nature of the protagonist-adversary learning paradigm, their parameter update principle can be uniformly formalized as follows: At timestep $t$, the protagonist executes action $a_{t}=a_{p_{t}}$ and receives reward $R_{t}=r_{t}$, while the adversary selects action $a_{t}=a_{a_{t}}$ and obtains reward $R_{t}=-r_{t}$. To reduce the variance in the Q-function estimate, we update parameter $\theta_{k}$ for each critic $Q_{\theta_{k}}(k\in\{1,2,\dots,K\})$ with the same target to minimize the soft Bellman residual:
@@
where $\alpha$ is the temperature parameter, and $Q_{\bar{E}}(s_{t},a_{t})$ represents the ensemble Q-value estimate aggregated from $K$ target critics through our proposed TDU function, whose formal definition will be detailed in the next section. Then, the policy $\pi_{\phi}$ can be updated by ensemble Q-value estimate and minimizing:
@@
where $Q_{E}(s_{t},a_{t})$ is the TDU output aggregated from $K$ critics.
(a) Robustness evaluation curves under different methodological configurations.
Diversity Enhancement.
To encourage sufficient diversity among critics, we employ a dual randomization strategy combining parameter initialization and architectural design. Each critic network undergoes independent Gaussian weight initialization with parameter noise injection, establishing distinct initial optimization paths (Plappert et al., 2018). Additionally, we randomly assign activation functions (ReLU, LeakyReLU, ELU) (Dubey et al., 2022) to each layer during forward propagation. This combined parametric and structural randomization systematically promotes representational diversity across critics while maintaining functional compatibility, with experimental results in Figure LABEL:diversity empirically validating its effectiveness.
4.2 Time-varying Decay Uncertainty Mechanism
Conventional ensemble methods in RL typically stabilize training by averaging critic outputs, which has been proven effective in reducing variance of target approximation errors (Anschel et al., 2017) and refining policy optimization directions (Chen et al., 2021b). Building on this foundation, we propose the TDU function, formally defined in Definition 1, as an optimistic aggregation mechanism tailored for $K$-critic ensembles in non-stationary environments. The TDU mechanism explicitly considers epistemic uncertainty among critics through variance quantification and dynamically modulating its impact on Q-value estimation via time-decaying coefficients. This dual mechanism design enables TDU to simultaneously enhance Q-value estimation stability and maintain an adaptive exploration-exploitation balance throughout the training process.
@@ 4.2 Time-varying Decay Uncertainty Mechanism
As evident from the above definition, the efficacy of the TDU function is principally governed by its time-varying decay coefficient $\beta(n)$, which dynamically modulates how epistemic uncertainty contributes to Q-value aggregation. Figure LABEL:coefficient compares the robustness performance of our proposed natural exponential decay formulation for $\beta(n)$ (given in Eq. (9)) with three baseline uncertainty-weighting strategies on the MuJoCo Hopper-Hop task: (i) uncertainty-agnostic: $\beta(n)\equiv 0$, equivalent to standard averaging; (ii) constant optimistic weighting: $\beta(n)\equiv 1$, fixed maximum uncertainty weighting; (iii) linear decay: $\beta(n)$ decreasing linearly from $1$ to 0. The results demonstrate that time-decaying weighting strategies enable agents to learn significantly more robust policies compared with both uncertainty-agnostic and constant optimistic weighting approaches. This advantage stems from their ability to automatically transition from exploration-emphasizing behavior in early training stages to confidence-prioritizing predictions in later phases, thereby improving both estimation accuracy and learning efficiency. Furthermore, our proposed natural exponential decay strategy achieves superior robustness performance and faster convergence than linear decay through its monotonically decreasing derivative, allowing faster reduction of the agent’s uncertainty preference during early training while preserving essential exploration capacity for effective learning progression.
4.3 Theoretical Guarantees for TDU
@@
    $E\big[Q_{\theta_{k}}(s_{t},a_{t})\big]=Q^{\ast}(s_{t},a_{t}),\forall k\in\{1,\dots,K\};$ (10)
@@
    $\operatorname{Var}\big[Q_{\theta_{k}}(s_{t},a_{t})\big]\leq\sigma^{2},\forall k\in\{1,\dots,K\}.$ (11)
@@
$\lim_{N\rightarrow\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq\beta_{\min}\cdot\sigma.$ (12)
As $\beta_{\min}\to 0$, the ensemble estimator $Q_{E}(s_{t},a_{t})$ converges in probability to $Q^{*}(s_{t},a_{t})$:
$\hskip 45.00006ptQ_{E}(s_{t},a_{t})\overset{p}{\to}Q^{\ast}(s_{t},a_{t}).$ (13)
@@
UACER is built upon established robust adversarial RL frameworks, including RARL (Pinto et al., 2017), MixedNE-LD (Kamalaruban et al., 2020), and QARL (Reddi et al., 2024), and thus inherits their theoretical guarantee of convergence to a Nash equilibrium under adversarial interactions. This foundation ensures stable and well-founded learning dynamics even in the presence of perturbations and strategic opponents. Moreover, Theorem 1 formally establishes the asymptotic reliability of TDU’s Q-value estimation. It shows that as the number of training iterations grows sufficiently large and $\beta_{\min}$ approaches zero, the expected bias between TDU’s Q-value estimates and the true Q-values diminishes asymptotically. This result confirms TDU’s consistency as an aggregation method and demonstrates its theoretical capability to converge to the optimal value function given sufficient training. The detailed proof of Theorem 1 is provided in Appendix A.
5 Experimental Results
  • [leftmargin=2em]
  • Q1
    Can UACER successfully achieve more robust (see Figures 4, 5 and Table 2), stable (see Table 1), and efficient (see Figure 4) adversarial RL compared with existing methods?
  • Q2
    How crucial is each proposed component in UACER for improving RL performance (see Figure 6)?
  • Q3
    Does the optimistic design of TDU provide measurable benefits in robustness and convergence compared to conventional pessimistic value estimation under adversarial conditions (see Figure 7)?
  • Q4
    How does critic ensemble size $K$ affect the performance (see Figure 9)?
  • Q5
    Can UACER be universally applicable to other existing robust adversarial RL algorithms (see Figure 8)?
Environments.
We evaluate our proposed method on a diverse set of MuJoCo control problems (Todorov et al., 2012) from the DeepMind Control Suite (Tunyasuvunakool et al., 2020), including Cartpole-Swingup Sparse, Walker-Run, Hopper-Hop, and Cheetah-Run. Following established adversarial training protocols (Pinto et al., 2017; Kamalaruban et al., 2020; Reddi et al., 2024), we instantiate two agents: a protagonist, which learns an optimal policy to accomplish the task, and an adversary, which exerts adversarial forces to disrupt the protagonist’s performance, thereby eliciting robust agent behavior. Both agents observe the standard environmental states as inputs. The adversary’s action space consists of bounded, multi-dimensional disturbance forces. For each environment, the adversary’s perturbation intensity is carefully tuned so as to be large enough to beget agent robustness and generalization while posing a challenge to the protagonist. Additional environment-specific parameter details are provided in Appendix B.1.
Comparison Methods.
We evaluate our approach in comparison to the classical robust adversarial RL method RARL (Pinto et al., 2017) and several recently proposed improvements addressing RARL’s saddle point optimization challenges, including MixedNE-LD (Kamalaruban et al., 2020), which utilizes Langevin dynamics to escape local optima, and Curriculum Adversarial Training (CAT) (Sheng et al., 2022) featuring dynamic adversary strength adjustment via a hand-designed curriculum. Additionally, we compare against the state-of-the-art QARL algorithm (Reddi et al., 2024), which synergistically combines entropy regularization with quantal response equilibrium while employing temperature-based curriculum learning to regulate adversary rationality. For comprehensive assessment, standard SAC (Haarnoja et al., 2018) serves as a non-adversarial baseline representing the underlying RL framework. Unless specified otherwise, UACER implementations in the following build upon the QARL framework. Complete hyperparameter configurations for all algorithms are provided in Appendix B.2.
Evaluation Protocols.
For each comparison method, we conduct training over five random seeds $\{0,1,2,3,4\}$ across all environments, with $200$ alternating iterations per seed. All reported results represent the average performance over these five runs. In the performance curves, each line depicts the average return obtained for varying properties of the environment after the corresponding training iteration. Solid lines indicate the mean performance, and shaded regions correspond to the $95\%$ confidence intervals across the five seeds. To evaluate the disturbance resistance of the trained protagonists to varying test conditions during training, we perform robustness evaluation every five alternating iterations. Specifically, for each environment, we assess the protagonist’s returns by sweeping through a predefined range of environmental parameter variations (e.g., mass and friction coefficient) in the absence of the adversary, and report the average return across all parameter configurations as the robustness metric. This procedure introduces graded disturbances through varying dynamics parameters, thereby offering a comprehensive test of the agent robustness while simultaneously assessing its generalization capability under domain shift conditions. Concurrently, we quantify training stability by computing the percentage decrease in the protagonist’s performance between consecutive (i.e., $i$ and $i+1$) robustness evaluations, then averaging these percentage decreases throughout the entire training process. In addition, following the practice in (Reddi et al., 2024), we assess the final adversarial performance of the protagonist in a minimax sense by fixing the protagonist obtained at the end of training and subsequently training an adversary against it, which is then used at test time.
Figure 5: Heatmap of final robustness evaluation for UACER and baseline methods on MuJoCo swingup and locomotion tasks.
5.1 Comparative Evaluation (Q1)
Robustness and Convergence.
As shown by the robustness evaluation results in Figure 4, UACER exhibits substantially accelerated convergence relative to all baseline methods across most environments, effectively addressing the persistent challenge of slow convergence in robust adversarial RL. This framework consistently achieves superior robustness scores across all environments, demonstrating both exceptional adversarial resilience and generalization capacity. Particularly in complex (Hopper-Hop) and sparse reward (Cartpole-Swingup Sparse) settings, UACER’s time-varying decay uncertainty-aware mechanism supports more efficient exploration and delivers statistically significant performance gains over competing approaches. Moreover, we visualize the robustness of the trained protagonist to different test conditions upon training completion, as illustrated in Figure 5. Each heatmap depicts the performance obtained under varying environmental parameters represented along the x- and y-xes. These results further validate the comprehensive robustness advantage of UACER under graded environmental perturbations.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
SAC $60.504\pm 53.811$ $353\pm 1.862 33.$$3.367\pm 8.704$ $4.520\pm 4.724$
CAT $56.902\pm 51.160$ $4.743\pm 2.955$ $32.276\pm 12.829$ $5.353\pm 3.882$
QARL $35.714\pm 28.414$ $4.527\pm 0.847$ $35.581\pm 6.025$ $8.692\pm 4.927$
UACER (ours) $17.352\pm 9.593$ $1.659\pm 0.604$ $27.938\pm 8.474$ $7.665\pm 3.515$
Table 1: Training stability analysis (percentage decrease, $\%$) of UACER and baseline methods on MuJoCo swingup and locomotion tasks.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
UACER (ours) $334.770\pm 6.886$ $341.995\pm 8.480$ $96.564\pm 17.479$ $312.594\pm 30.422$
Table 2: Final adversarial performance of UACER and baseline methods on MuJoCo swingup and locomotion tasks.
Training Stability.
We evaluate the training stability of UACER from two aspects: variability across different random seeds and consistency of returns during training. Figure 4 shows UACER achieves the smallest variance across seeds (narrowest shaded regions), indicating the lowest sensitivity to initialization for training. Furthermore, training stability metrics derived from the results in Figure 4 are summarized in Table 1, where performance (percentage decrease, $\%$) is reported as mean $\pm$ standard error across five independent trials, with bold and underlined values denoting the best and second-best results, respectively. The results show that UACER achieves the best training stability in the first three environments while maintaining comparable performance to the baselines in the last environment (i.e., Cheetah-Run). These results further validate its effectiveness in mitigating adversarial-induced training fluctuations in robust adversarial RL settings.
Worst-Case Performance.
As quantified final adversarial performance in Table 2, UACER consistently maintains superior performance with a relatively small variance across all environments when confronting optimized adversaries. This validates its dual capability of preserving policy effectiveness under distributional shifts while enhancing generalizability in extreme out-of-distribution scenarios.
5.2 Ablation Study (Q2)
We systematically evaluate the individual contributions of UACER’s core components: the critic ensemble, the TDU aggregation mechanism, and the diversity enhancement strategies. Through controlled single-variable ablation experiments, we compare three algorithmic variants in adversarial RL settings: (i) w/o ensemble: Replaces the $K$-critic ensemble with SAC’s standard two-critic architecture; (ii) w/o TDU: Substitutes TDU aggregation function with minimum-value selection used in SAC; (iii) w/o diversity: Removes the diversity enhancement strategies from UACER.
The experimental results in Figure 6 demonstrate that removing any component from UACER degrades performance substantially. When eliminating the diversity enhancement strategies ("w/o diversity"), the algorithm exhibits markedly poorer training stability and robustness across most environments. More critically, the removal of either the TDU mechanism ("w/o TDU") or the critic ensemble ("w/o ensemble") leads to notable performance declines in both robustness and convergence speed, particularly in challenging environments such as Cartpole-Swingup Sparse and Hopper-Hop. In contrast, the fully integrated UACER framework consistently outperforms all ablated versions. These results indicate that under highly perturbed environments, no individual component suffices for robust performance; instead, their synergistic integration is essential.
5.3 Superiority of TDU’s Optimistic Aggregation in Adversarial RL (Q3)
[…]
To evaluate the superiority of TDU’s optimistic aggregation in adversarial RL, we compare it with three pessimistic variants derived from Eq. (6): (i) pessimism-dec: the mean minus the uncertainty term, with $\beta(n)$ decaying from 1 to 0; (ii) pessimism-inc: the mean minus the uncertainty term, with $\beta(n)$ increasing from 0 to 1; (iii) pessimism-min: the minimum among the $K$ critic outputs. As demonstrated in Figure 7, TDU consistently delivers superior performance across all tasks. This advantage arises because, in non-stationary adversarial environments, the protagonist already struggles to learn appropriate responses to perturbations, and pessimistic estimation further constraints its Q-value prediction capacity under such disturbances. By contrast, TDU’s optimistic aggregation achieves a more effective balance between bias control and robustness enhancement, thereby alleviating the tendency toward over-conservative policy learning while simultaneously strengthening training stability and final performance.
Figure 7: Robustness evaluation curves of UACER (optimistic TDU aggregation) and its three pessimistic Q-value estimation variants on MuJoCo swingup and locomotion tasks.
Figure 8: Robustness evaluation curves of UACER combined with RARL/SAC MixedNE-LD on MuJoCo swingup and locomotion tasks.
Figure 9: Robustness evaluation of UACER with varying values of critic ensemble size $K$ on the Hopper-Hop environment.
5.4 Hyperparameter and Scalability Analysis
Effects of Critic Ensemble Size (Q4)
The ensemble size $K$ (number of critics) in UACER critically influences the agent’s performance. To investigating this relationship, we conduct controlled experiments varying $K$ from $2$ to $10$ while holding other hyperparameters fixed. As illustrated in Figure 9, we observe that the final obtained return in terms of robustness evaluation improves gradually with increasing $K$ up to $K=5$, after which performance declines. This indicates that while larger ensembles enhance uncertainty estimation and ensemble diversity, excessive critics ($K>5$) may introduce detrimental redundancy, complicating policy optimization. Consequently, a higher value of $K$ is not always better. Here, for the Hopper-Hop task, $K=5$ achieves the optimal robustness gains, suggesting task-independent determination of $K$ is essential.
Scalability Analysis (Q5)
All previous experiments implemented UACER based on the QARL framework as the foundational robust adversarial RL approach. This section examines UACER’s applicability as a general framework when integrated with other existing robust adversarial RL algorithms. Specifically, we migrate UACER to the classic RARL algorithm and its improved extension SAC MixedNE-LD for practical implementation and evaluation. The learning curves in terms of robustness performance are shown in Figure 8. The experimental results confirm that both RARL and SAC MixedNE-LD achieve substantial performance gains after incorporating UACER, showing improved convergence speed, training stability, and robustness to adversarial perturbations. In essence, UACER is designed as an algorithm-agnostic framework that is compatible with any value-based robust adversarial RL methods. Regrettably, in the Cartpole-Swingup Sparse task, SAC MixedNE-LD exhibited critically poor baseline performance (near-zero returns throughout training), with no discernible improvement after UACER integration. This outcome indicates that UACER’s performance improvements remain partly dependent on the base adversarial RL algorithm’s intrinsic adaptability to the specific task.
5.5 Computational Overhead and Limitations
6 Conclusion
We propose UACER, a novel approach that significantly advances robust adversarial RL methods through two key contributions: (i) a parallel ensemble of critic networks that simultaneously reduces Q-value estimation variance and enhances robustness against adversarial perturbations, and (ii) a Time-varying Decay Uncertainty (TDU) Q-value aggregation mechanism that dynamically balances exploration and exploitation through variance-based epistemic uncertainty quantification across critics. Theoretically grounded and extensively validated, UACER demonstrates substantial improvements in training stability, convergence speed, and adversarial robustness, addressing critical challenges for real-world deployment of deep RL systems. By integrating ensemble learning with uncertainty-aware optimization, our framework enables the development of adversarial RL algorithms with proven stability guarantees and robust performance in complex and uncertain environments. Future research directions include extending UACER to partially observable environments and investing its applications in real-world safety-critical systems such as robotic manipulation and autonomous driving scenarios.
Appendix A Proof of Theorem 1
In this section, we present the proof of Theorem 1. Using the conditions specified in Definition 1, we establish the theorem’s validity as follows.

### v2

@@ IV Methodology
In this section, we give a detailed description of UACER, whose overview is illustrated in Fig. 1. We first introduce the diversified critic ensemble framework for robust adversarial RL and then propose the variance-based TDU Q-value aggregation function, followed by theoretical guarantees for TDU’s convergence properties and aggregation effectiveness. In principle, our method can be used in conjunction with most modern robust adversarial RL algorithms, including recent advances like QARL [30] as well as classical approaches such as RARL [28]. For the sake of concreteness, we here present the QARL-based implementation as our primary instantiation.
IV-A Diversified Critic Ensemble
The Unified Framework
@@ IV-A Diversified Critic Ensemble
The UACER framework implements both protagonist and adversary as actor-critic agents based on the SAC algorithm. Each agent consists of a single policy network, denoted as $\pi_{\phi}$, and an ensemble of $K$ soft Q-function networks $\{Q_{\theta_{k}}\}_{k=1}^{K}$. This design is motivated by the empirical insight, as shown in Fig. 2(a), that the critic ensemble significantly enhances the robustness of Q-value estimation, especially under adversarial perturbations. By evaluating Q-values through multiple independent critics, the framework mitigates the influence of localized noise or perturbations on any single estimator, thereby yielding more stable and reliable value estimates. In contrast, although actor ensembles can improve sample efficiency and generalization [20, 39], they often introduce additional training instability due to policy divergence among actors, which may produce conflicting behaviors that undermine the consistency and convergence of the learned policy.
Algorithm 1 UACER
0: Initial $K$ critics for protagonist: $\{Q_{\theta_{k}^{p}}\}_{k=1}^{K}$, $\{Q_{\bar{\theta}_{k}^{p}}\}_{k=1}^{K}$;  Initial $K$ critics for adversary: $\{Q_{\theta_{k}^{a}}\}_{k=1}^{K}$, $\{Q_{\bar{\theta}_{k}^{a}}\}_{k=1}^{K}$;  Initial actor $\pi_{\phi_{p}}$, $\pi_{\phi_{a}}$ for protagonist and adversary;  Replay buffer $\mathcal{D}_{p}=\emptyset$, $\mathcal{D}_{a}=\emptyset$;  Total number of iterations NN;  TDU related parameters $\beta_{0},\beta_{\min},\lambda$.
0: Protagonist’s policy $\pi_{\phi_{p}}$.
1:for each iteration $n\in\{1,\dots,N\}$ do
2:  Update target critics for both agents:  $\{\bar{\theta}_{k}^{p}\}_{k=1}^{K}\leftarrow\{\theta_{k}^{p}\}_{k=1}^{K}$, $\{\bar{\theta}_{k}^{a}\}_{k=1}^{K}\leftarrow\{\theta_{k}^{a}\}_{k=1}^{K}$.
3:  while performing the adversary update do
4:   Both agents interact with the environment and store the collected transitions in $\mathcal{D}_{a}$.
5:   Update $\{Q_{\theta_{k}^{a}}\}_{k=1}^{K}$ using Eqs.(5), (6) and (8).
6:   Update $\pi_{\phi_{a}}$ using Eqs.(7) and (8).
7:  end while
8:  while performing the protagonist update do
9:   Both agents interact with the environment and store the collected transitions in $\mathcal{D}_{p}$.
10:   Update $\{Q_{\theta_{k}^{p}}\}_{k=1}^{K}$ using Eqs.(5), (6) and (8).
11:   Update $\pi_{\phi_{p}}$ using Eqs.(7) and (8).
12:  end while
13:end for
14:return $\pi_{\phi_{p}}$.
@@
During the adversarial training phase, both agents interact synchronously with the environment through their current policies, generating transition tuples $\tau_{t}=(s_{t},a_{p_{t}},a_{a_{t}},r_{t},s_{t+1})$ that are stored in the experience replay buffer $\mathcal{D}$. Given the symmetric nature of the protagonist-adversary learning paradigm, their parameter update principle can be uniformly formalized as follows: At timestep $t$, the protagonist executes action $a_{t}=a_{p_{t}}$ and receives reward $R_{t}=r_{t}$, while the adversary selects action $a_{t}=a_{a_{t}}$ and obtains reward $R_{t}=-r_{t}$. To reduce the variance in the Q-function estimate, we update parameter $\theta_{k}$ for each critic $Q_{\theta_{k}}(k\in\{1,2,\dots,K\})$ with the same target to minimize the soft Bellman residual:
@@
where $Q_{\bar{E}}(s_{t},a_{t})$ represents the ensemble Q-value estimate aggregated from $K$ target critics through our proposed TDU function, whose formal definition will be detailed in the next section. Then, the policy $\pi_{\phi}$ can be updated by the ensemble Q-value estimate and minimizing:
@@
where $Q_{E}(s_{t},a_{t})$ is the TDU output aggregated from $K$ critics. The complete iterative optimization procedure of our unified framework is summarized in Algorithm 1.
Figure 2: Robustness evaluation curves of (a) ensemble component ablation, (b) diversity ablation, and (c) different uncertainty coefficients on MuJoCo swingup and locomotion tasks. Here and in related figures in the following, each curve shows the average return obtained for varying properties of the environment after the corresponding training iteration. The solid lines and shaded regions denote the average performance and the $95\%$ confidence intervals, respectively, across five seeds.
Diversity Enhancement
To encourage sufficient diversity among critics, we employ a dual randomization strategy combining parameter initialization and architectural design. Each critic network undergoes independent Gaussian weight initialization with parameter noise injection, establishing distinct initial optimization paths [29]. Additionally, we randomly assign activation functions (ReLU, LeakyReLU, ELU) [10] to each layer during forward propagation. This combined parametric and structural randomization systematically promotes representational diversity across critics while maintaining functional compatibility, with experimental results in Fig. 2(b) empirically validating its effectiveness.
IV-B Time-varying Decay Uncertainty Mechanism
Conventional ensemble methods in RL typically stabilize training by averaging critic outputs, which has been proven effective in reducing variance of target approximation errors [3] and refining policy optimization directions [7]. Building on this foundation, we propose the TDU function, formally defined in Definition 1, as an optimistic aggregation mechanism tailored for $K$-critic ensembles in non-stationary environments. The TDU mechanism explicitly considers epistemic uncertainty among critics through variance quantification and adaptively modulates its impact on Q-value estimation via time-decaying coefficients. This dual mechanism design enables TDU to simultaneously enhance Q-value estimation stability and maintain an adaptive exploration-exploitation balance throughout the training process.
@@ IV-B Time-varying Decay Uncertainty Mechanism
As evident from the above definition, the efficacy of the TDU function is principally governed by its time-varying decay coefficient $\beta(n)$, which adaptively modulates how epistemic uncertainty contributes to Q-value aggregation. Fig. 2(c) compares the robustness performance of our proposed natural exponential decay formulation for $\beta(n)$ (given in Eq. (11)) with three baseline uncertainty-weighting strategies on the MuJoCo Hopper-Hop task: (i) uncertainty-agnostic: $\beta(n)\equiv 0$, equivalent to standard averaging; (ii) constant optimistic weighting: $\beta(n)\equiv 1$, fixed maximum uncertainty weighting; (iii) linear decay: $\beta(n)$ decreasing linearly from $1$ to 0. The results demonstrate that time-decaying weighting strategies enable agents to learn significantly more robust policies compared with both uncertainty-agnostic and constant optimistic weighting approaches. This advantage stems from their ability to adaptively transition from exploration-emphasizing behavior in early training stages to confidence-prioritizing predictions in later phases, thereby improving both estimation accuracy and learning efficiency. Furthermore, our proposed natural exponential decay strategy achieves superior robustness performance and faster convergence than linear decay through its monotonically decreasing derivative, allowing faster reduction of the agent’s uncertainty preference during early training while preserving essential exploration capacity for effective learning progression.
IV-C Theoretical Guarantees for TDU
@@
    $E[Q_{\theta_{k}}(s_{t},a_{t})]=Q^{\ast}(s_{t},a_{t}),\forall k\in\{1,\dots,K\};$ (12)
@@
    $\operatorname{Var}[Q_{\theta_{k}}(s_{t},a_{t})]\leq\sigma^{2},\forall k\in\{1,\dots,K\}.$ (13)
@@
$\lim_{N\rightarrow\infty}\mathbb{E}\big[|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})|\big]\leq\beta_{\min}\cdot\sigma.$ (14)
As $\beta_{\min}\to 0$, the ensemble estimator $Q_{E}(s_{t},a_{t})$ converges in probability to $Q^{\ast}(s_{t},a_{t})$:
$Q_{E}(s_{t},a_{t})\overset{p}{\to}Q^{\ast}(s_{t},a_{t}).$ (15)
@@
UACER is built upon established robust adversarial RL frameworks, including RARL [28], MixedNE-LD [17], and QARL [30], and thus inherits their theoretical guarantee of convergence to a Nash equilibrium under adversarial interactions. This foundation ensures stable and well-founded learning dynamics even in the presence of perturbations and strategic opponents. Moreover, Theorem 1 formally establishes the asymptotic reliability of TDU’s Q-value estimation. It shows that as the number of training iterations grows sufficiently large and $\beta_{\min}$ approaches zero, the expected bias between TDU’s Q-value estimates and the true Q-values diminishes asymptotically. This result confirms TDU’s consistency as an aggregation method and demonstrates its theoretical capability to converge to the optimal value function given sufficient training.
The proof of Theorem 1 is provided as follows.